OpenStudy (alyssakin):

If (7^2)^p = 7^6, what is the value of p? I am confused on this one. Do I still add the exponents so p would be 4 or because of the parenthesis do I multiply and the p would be 3?

OpenStudy (ibrafacts):

7

OpenStudy (ibrafacts):

7^2 = 49

OpenStudy (anonymous):

\[7^{2p} = 7^6.\] So, 2p = 6.

OpenStudy (anonymous):

Actually is multipled to the 2 so p=3

OpenStudy (ibrafacts):

yes my bad didnt see p is ^ thought it was an add sign sorry

OpenStudy (alyssakin):

So because of the parenthesis in this case we do multiply the exponents of 2*3 to equal 6? and not add them as 2+4=6? Just to clarify. Thank you

OpenStudy (anonymous):

Yeah because it is in parentheses and there is an exponent outside you multiply, now if you were to have 7^2*7^3, then you would add and get 7^5

OpenStudy (alyssakin):

Ok thank you!!!